# Leave-Group-Out Cross-Validation for Latent Gaussian Models

### v5

@@ Abstract
Evaluating the predictive performance of a statistical model is commonly done using cross-validation. Although the leave-one-out method is frequently employed, its application is justified primarily for independent and identically distributed observations. However, this method tends to mimic interpolation rather than prediction when dealing with dependent observations.
This paper proposes a modified cross-validation for dependent observations. This is achieved by excluding an automatically determined set of observations from the training set to mimic a more reasonable prediction scenario. Also, within the framework of latent Gaussian models, we illustrate a method to adjust the joint posterior for this modified cross-validation to avoid model refitting. This new approach is accessible in the R-INLA package (www.r-inla.org).
@@ 1 Introduction
1.1 Background
Leave-one-out cross-validation (LOOCV) [1] stands as a popular method for evaluating a statistical model’s predictive performance, model selections, or estimating some critical parameters in the model. The core concept of LOOCV is elegantly straightforward. Suppose we have data, $\boldsymbol{y}=\{y_{i}\}$, for $i=1,\ldots,n$, presumed to be Independent and Identically Distributed (I.I.D.) samples from the true distribution $\pi_{T}(y)$. Our objective is to determine how well a fitted model can predict a new observation, $\tilde{y}$, sampled from this true distribution. In the Bayesian context, we use the posterior predictive distribution $\pi(y|\boldsymbol{y})$ to predict $\tilde{y}$ sampled from $\pi_{T}(y)$. By using the logarithmic score [2], we compute $E_{\tilde{y}}[\log\pi(\tilde{y}|\boldsymbol{y})]$ as a metric for prediction quality.
Owing to the lack of $\pi_{T}(y)$, directly computing the expectation becomes infeasible. Nonetheless, since $y_{i}$ is an I.I.D. sample of $\pi_{T}(y)$, we can estimate this expectation by evaluating
@@ 1.1 Background
However, issues arise when the I.I.D. assumption does not hold. We can have longitudinal data following each subject in a study [3], dependence due to time and/or space [4, 5], or hierarchical structure [6]. In those cases, LOOCV often has an overly optimistic assessment of models’ predictive performance since LOOCV no longer mimics a proper prediction task, which defines the new data generation process given observed data. This paper proposes an adaptation of LOOCV that is more aligned with “measuring predictive performance” when the I.I.D. assumption is invalid.
@@ 1.2 The prediction task
The critical observation is that the meaning of “prediction" is not clearly defined when $\boldsymbol{y}$ are not I.I.D samples of $\pi_{T}(y)$. $\pi_{T}(\tilde{y}|\boldsymbol{y})$ lacks a unique definition in non-I.I.D. scenarios as without a clear prediction task, i.e., how we imagine a new future data point, $\tilde{y}$, is generated given observed data $\boldsymbol{y}$. This ambiguity extends to the act of “using $\boldsymbol{y}$ to predict $\tilde{y}$" as it is uncertain what our target, $\tilde{y}$, represents. To illustrate these concepts, let us discuss some more concrete examples.
@@
Assume data $\boldsymbol{y}=\{y_{1},y_{2},\dots,y_{T}\}$ is a time-series, observed sequentially at time $1,2,\dots,T$. The inherent prediction task is to predict future values, given the temporal nature of the data. We can predict a new observation at $k\geq 1$ steps ahead into the future by $\pi(y_{T+k}|y_{1},\dots,y_{T})$.
@@
which is often referred as interpolation or imputation of missing values rather than a prediction. However, time series models’ predictive performance is often assessed through leave-future-out cross-validation (LFOCV) [5]:
@@
The message from this example is that LOOCV, when applied to such models, is essentially evaluating interpolation performance rather than predictive performance.
We acknowledge two issues. First, the distinction between interpolation and prediction is not always clear-cut, leading to overlapping concepts. For example, a one-step-ahead forecast leans more towards interpolation than a two-step-ahead prediction. In contrast, a one-step-ahead forecast leans less towards interpolation than a missing value imputation. However, this does not deter our discussion. Secondly, while an ideal model succeeds in all prediction tasks, real-world scenarios demand us to settle for the definition of the “best fit". Consequently, our choice of evaluation should align with our specific objectives.
@@
Figure 1 illustrates an example of a multilevel model. Consider observations of student grades or performance. This data exhibits a hierarchical structure: students belong to classes, classes reside within schools, and schools are nested within regions. This hierarchical arrangement is significant because it introduces effects attributed to the class, school, and region levels, substantially deviating from the I.I.D. assumption.
@@
To properly evaluate predictive performance within this context, users must first explicitly define their prediction task and then assess the model in line with this definition. It is worth noting that applying LOOCV would imply the prediction of individual students within observed classes. In our view, this mimics more interpolation rather than prediction.
@@
1.3 Improving LOOCV for non-I.I.D. models
Our discussions illuminate an important insight: when dealing with non-I.I.D. data, the prediction task implicitly defined through LOOCV may be less appropriate, as it leans more towards assessing interpolation qualities than predictive performance. This prompts the question: What is a suitable approach moving forward?
@@
Here, the group (denoted by $I_{i}$) is an index set including $i$. This configuration facilitates that the pair $(y_{i},\boldsymbol{y}_{-I_{i}})$ mimics a specified prediction task, with $\boldsymbol{y}_{-I_{i}}$ being the data subset excluding the data indexed by $I_{i}$. In a multilevel model, as depicted in Figure 1, predicting a student’s grade from an unseen class necessitates that $I_{i}$ includes $i$ and all observations from student $i$’s class. However, more complex models, such as models containing both time series and hierarchical elements, pose challenges when defining a natural prediction task. Therefore, unless in simple cases otherwise, LOOCV is often applied for its simplicity—even if it leans more towards interpolation.
Our primary contribution is offering an automatic LGOCV adaptive to the specified model when prediction tasks are not defined. Our approach automatically constructs a group, $I_{i}$, for each $i$. Though we will delve into automatically defining $I_{i}$ later, an initial understanding is that $I_{i}$ comprises the data points most informative for predicting the testing point, $y_{i}$. This set ensures that our LGOCV focuses less on interpolation and more on prediction than LOOCV. In other words, LGOCV tests the model on more difficult prediction tasks.
The rationale behind LGOCV is to create more reasonable criteria than LOOCV when the user has difficulty defining manual groups in non-I.I.D. models. In various practical examples, we will show how this automatic procedure produces reasonable groups. Also, many advanced spatial examples applying this method are shown in the published paper [7]. For a simple time-series example, our new approach will correspond to evaluating $\pi(y_{t}|y_{1},\ldots,y_{t-k},y_{t+k},\ldots,y_{T}),$ for fixed $k>1$. This corresponds to removing a sequence of data with length $12k-$ 1, to predict the central one. As we see, this is neither pure interpolation nor pure prediction. Our interpretation is that it is less interpolation or more prediction than what LOOCV provides when $k>1$.
@@
There are two key challenges to address to make our proposal practical. Firstly, we must quantify the information contributed by one data point in predicting another; this is crucial for automatic group construction. Secondly, we face the computational task of evaluating $u_{\text{LGOCV}}$ given a set of groups. The naive computation of LGOCV by fitting models across all potential training sets and evaluating their utility against corresponding testing points is computationally infeasible, especially given the resource-demanding nature of modern statistical models. However, these challenges can be handled elegantly within the framework of latent Gaussian models (LGMs) combined with the integrated nested Laplace approximation (INLA) inference, as detailed in [8, 9, 10, 11]. Throughout this paper, we will assume that our model is an LGM. We will discuss how to integrate the automatic group construction and the fast computation of $u_{\text{LGOCV}}$ into the LGMs-INLA framework. Notably, our proposed methodology has been incorporated into the R-INLA package (www.rinla.org), extending its applicability across all LGMs supported by R-INLA.
1.4 Related studies
The issues of applying LOOCV to non-I.I.D. models have been recognized and addressed by numerous researchers. To name a few, [4] advocate block cross-validation, partitioning ecological data based on inherent patterns, but this approach should be adopted when the prediction task is not simply interpolation. [12] offers a modification to LOOCV, ensuring an unbiased measure of predictive performance given the correlation between new and observed data, where the unbiasedness is in the sense of randomized both observed data and new data. We should note that an assumed prediction task determines the correlation between new and observed data. [5] presents an efficient approximation for LFOCV, which is used when the prediction task is forecasting in time series. [13] considers a multilevel model and demonstrates that marginal WAIC is akin to LOOCV. In contrast, conditional WAIC aligns with LGOCV, where a hierarchical level, such as a school, defines the groups. The choice of criteria is according to one’s assumption of the new data, i.e., the prediction task. Additionally, [14] recommends h-block cross-validation to estimate one-step-ahead performance in stationary processes, which is similar to applying our approach to a stationary model.
These studies assume that model users can select an appropriate evaluation method judiciously according to their understanding of their prediction tasks. Yet, in practice, many users, despite knowing LOOCV’s limitations, still resort to LOOCV or randomized K-fold cross-validation, primarily for user-friendliness. This reality motivates our pursuit of a method that provides user-friendliness with better predictive performance estimation. To this extent, an R package blockCV [15] implements spatial and environmental blocking and spatial buffering. Spatial blocking forms clusters of data points according to spatial effects, and environmental blocking forms clusters using K-means [16] on the covariates. Buffering is the same as our cross-validation with the groups formed only by spatial effects. This approach ensures that no test data is spatially next to any training data. This package is written mainly for species distribution modeling, while our approach is designed for a more general purpose. If we put our method in the context of species distribution modeling, our method provides buffers that are generated by a posterior Gaussian process considering all the relevant model effects, including spatial, temporal, and environmental factors. In other words, our method ensures that no testing data abuts training data in terms of the combination of spatial and environmental effects.
@@ 1.5 Theoretical aspects
Cross-validation (CV), particularly LOOCV, is frequently considered as an estimator of $E_{\tilde{y}}[\log\pi(\tilde{y}|\boldsymbol{y})]$ or $E_{\tilde{y},\boldsymbol{y}}[\log\pi(\tilde{y}|\boldsymbol{y})]$. The first expectation describes the generalized predictive performance given a specific training set, while the second expectation describes the generalized predictive performance averaged over different identically distributed training sets. These expectations can be evaluated when assuming the existence of the joint density $\pi_{T}(\tilde{y},\boldsymbol{y})$, representing the true data generation process. Under the assumption of I.I.D. and some regularity conditions on the model, the Bernstein-Von-Mises theorem states that $\log\pi(\tilde{y}|\boldsymbol{y})$ converges to a random variable irrelevant to $\boldsymbol{y}$. Consequently, $E_{\tilde{y}}[\log\pi(\tilde{y}|\boldsymbol{y})]$ and $E_{\tilde{y},\boldsymbol{y}}[\log\pi(\tilde{y}|\boldsymbol{y})]$ become equivalent in the limit. If we further assume that $\tilde{y}$ is sampled from the same distribution as all the training data, LOOCV is an asymptotically unbiased estimator of the expectations. Commonly used information criteria, such as AIC [17], WAIC [18], are asymptotically equivalent to LOOCV in I.I.D. cases.This type of analysis is prevalent in the literature with various settings [1, 19, 20, 21].
However, a similar analysis does not hold in non-I.I.D. cases in general. Firstly, the existence of different prediction tasks means that both the model prediction, $\pi(\tilde{y}|\boldsymbol{y})$, and the true data generation process, $\pi_{T}(\tilde{y}|\boldsymbol{y})$, are not uniquely defined as discussed in section 1.2. Secondly, the asymptotic scheme is not uniquely defined, even with a specific prediction task. For example, in a temporal model where data $\boldsymbol{y}=\{y_{1},y_{2},\dots,y_{n}\}$ is a time-series, observed at time $t_{1}<t_{2}<\dots<t_{n}$ and we denote the last time step as $T$. Several meanings of of $n\to\infty$ can be considered:
@@
These scenarios correspond to observing more future data and having higher sample rates within a time frame. As mentioned in the section 1.2, multilevel data can also have various asymptotic regimes. Thirdly, if the data generation process is not stationary, the model will not converge under certain asymptotic regimes, which differentiate $E_{\tilde{y}}[\log\pi(\tilde{y}|\boldsymbol{y})]$ from $E_{\tilde{y},\boldsymbol{y}}[\log\pi(\tilde{y}|\boldsymbol{y})]$ even in asymptotic scenarios. These points highlight that the estimand of CV is not uniquely defined in non-I.I.D. cases, preventing the establishment of asymptotic analysis frameworks similar to I.I.D. analysis.
From the perspective of CV, it is also inappropriate to consider it an estimator since each summand in CV should be viewed as a sample from different distributions due to the relevance of data indexes in non-I.I.D. scenarios. For example, if we compute LOOCV in a time series. Each $y_{t}$ is sampled from different conditional distribution $\pi_{T}(y_{t}|\boldsymbol{y}_{-t})$ and thus the average $\frac{1}{n}\sum_{t=1}^{T}\log\pi(y_{t}|\boldsymbol{y}_{-t})$ can not be considered as an estimator in general. Therefore, it is more reasonable to view CV as a predictive measurement rather than an estimator of an expectation. This perspective allows us to interpret the proposed automatic LGOCV as averaged predictive performance for similar prediction tasks created automatically by the model.

### v6

@@ Abstract
Evaluating the predictive performance of a statistical model is commonly done using cross-validation. Among the various methods, leave-one-out cross-validation (LOOCV) is frequently used. Originally designed for exchangeable observations, LOOCV has since been extended to other cases such as hierarchical models. However, it focuses primarily on short-range prediction and may not fully capture long-range prediction scenarios. For structured hierarchical models, particularly those involving multiple random effects, the concepts of short- and long-range predictions become less clear, which can complicate the interpretation of LOOCV results. In this paper, we propose a complementary cross-validation framework specifically tailored for longer-range prediction in latent Gaussian models, including those with structured random effects. Our approach differs from LOOCV by excluding a carefully constructed set from the training set, which better emulates longer-range prediction conditions. Furthermore, we achieve computational efficiency by adjusting the full joint posterior for this modified cross-validation, thus eliminating the need for model refitting. This method is implemented in the R-INLA package (www.r-inla.org) and can be adapted to a variety of inferential frameworks.
@@ 1 Introduction
1.1 Rationale and Background
Leave-one-out cross-validation (LOOCV) [1] stands as a popular method to evaluate a statistical model’s predictive performance, perform model selections, or estimating some critical parameters in the model. The core concept of LOOCV is elegantly straightforward. Suppose we have data, $\boldsymbol{y}=\{y_{i}\}$, for $i=1,\ldots,n$, presumed to be Independent and Identically Distributed (I.I.D.) samples from the true distribution $\pi_{T}(y)$. Our objective is to determine how well a fitted model can predict a new observation, $\tilde{y}$, sampled from this true distribution. In the Bayesian context, we use the posterior predictive distribution $\pi(y|\boldsymbol{y})$ to predict $\tilde{y}$ sampled from $\pi_{T}(y)$ as proposed by [2]. Using the logarithmic score [3], we can compute $E_{\tilde{y}}[\log\pi(\tilde{y}|\boldsymbol{y})]$ as a metric for prediction ability.
Owing to the lack of $\pi_{T}(y)$, directly computing the expectation becomes infeasible. Nonetheless, since $y_{i}$ is an exchangeable sample from $\pi_{T}(y)$, we can estimate this expectation by evaluating
@@ 1.1 Rationale and Background
However, issues can arise in more complex statistical models where the dependency in the model results in the data not being exchangeable (see [4] for a complete discussion of CV for several types of exchangeability); we describe these kinds of models as "dependent" cases for the purpose of this paper. An intuitive dependent case is a time series. [5] proposed a block CV method for dependent data from a stationary process, acknowledging the need for a different approach to CV than LOOCV. [6] propose modified cross-validation (MCV) where dependent data chunks are removed together with the relevant point to account for the dependence in a time series (and other dependent data generating models). [7] investigated the properties of blocked CV and other approaches for robust time series model evaluations (see also [8] for a study on k-fold CV), while [9] proposed a leave-future-out CV strategy. [10] investigated CV and holdout approaches for time series models and concluded that the out-of-sample holdout procedure is more accurate for non-stationary processes than LOOCV.
Besides time series, spatial dependence models come to mind for which [11] proposed a buffering strategy by leaving out specific spatial points or areas and spatial and environmental blocking. Spatial blocking forms clusters of data points according to spatial effects, and environmental blocking forms clusters using K-means [12] on the covariates. Other examples of dependent cases are longitudinal data for multiple subjects in a study [13] and hierarchical models (see [14] and [4, Section 5.1.4.]). [15] proposed an hv-block CV approach for dependent data while [16] considers a multilevel model and shows that marginal WAIC is akin to LOOCV. [17] advocate a block cross-validation, partitioning ecological data based on inherent patterns, when the prediction task is not simply short-range prediction. [18] offers a modification to LOOCV, ensuring an unbiased measure of predictive performance given the correlation between new and observed data, where the unbiasedness is in the sense of randomized both observed and new data. We should note that an assumed prediction task determines the correlation between new and observed data.
In dependent cases, LOOCV can provide a restricted assessment of the models’ predictive performance since LOOCV cannot evaluate longer-range prediction. Even in terms of short-range prediction, it is not clear what is short- or longer-range in dependent models that are not purely temporal or spatial models where the range has a physical interpretation. We use the concepts of short-range and longer-range predictions, acknowledging that these concepts can have overlapping meanings.
We thus propose a framework that emulates longer-range prediction scenarios, for hierarchical models, by constructing non-random leave-out sets based on model-based correlations. This can be viewed as a complementary approach to LOOCV for evaluating predictive performance, providing additional informative insights of the predictive ability for dependent cases.
@@ 1.2 The prediction task
The critical observation is that the meaning of “prediction" is not clearly defined when we are far away from exchangeability, so that $\boldsymbol{y}$ are non-exchangeable samples of $\pi_{T}(y)$. $\pi_{T}(\tilde{y}|\boldsymbol{y})$ lacks a unique definition in dependent cases as without a clear prediction task, i.e., how we imagine a new data point, $\tilde{y}$, is generated given observed data $\boldsymbol{y}$. This ambiguity extends to the act of “using $\boldsymbol{y}$ to predict $\tilde{y}$" as it is uncertain what our target, $\tilde{y}$, represents. To illustrate these concepts, let us discuss some more concrete examples.
@@
Assume data $\boldsymbol{y}=\{y_{1},y_{2},\dots,y_{T}\}$ is a time-series, observed sequentially at time $1,2,\dots,T$. The inherent prediction task is to predict future values, given the temporal nature of the data. We can predict a new observation at $k\geq 1$ steps into the future by $\pi(y_{T+k}|y_{1},\dots,y_{T})$.
@@
which is often referred to as interpolation or imputation of missing values, rather than a prediction. However, the predictive performance of time series models is often assessed through leave-future-out cross-validation (LFOCV) [9]:
@@
The message from this example is that LOOCV, when applied to such models, is essentially evaluating short-range prediction performance rather than longer-range predictive performance.
We acknowledge two issues. First, the distinction between short and longer-range prediction is not always clear-cut, leading to overlapping concepts. For example, a one-step-ahead forecast leans more towards short range than a two-step-ahead prediction. In contrast, a one-step-ahead forecast leans less towards short-range than a missing value imputation. However, this does not deter our discussion. Secondly, while an ideal model succeeds in all prediction tasks, real-world scenarios require us to settle for the definition of the “best fit". Consequently, our choice of evaluation should align with our specific objectives.
@@
Figure 1 illustrates an example of a multilevel model. Consider observations of student grades or performance. This data exhibits a hierarchical structure: students belong to classes, classes reside within schools, and schools are nested within regions. This hierarchical arrangement is significant because it introduces correlated random effects attributed to the class, school, and region levels, substantially deviating from the exchangeable case.
@@
To evaluate predictive performance within this context, users must first explicitly define their prediction task and then evaluate the model according to this definition. It should be noted that applying LOOCV would evaluate the prediction of individual students within observed classes. In our view, this mimics more short-range prediction rather than longer-range prediction, and another framework is needed to quantify the predictive ability for a new student in a new class in a new school in a new region, for example. Our proposal provides some insight into this kind of prediction task.
@@
1.3 LGOCV: Complementing LOOCV for dependent cases
Our discussions illuminate an important insight: when dealing with models that lead to non-exchangeable data, the prediction task implicitly defined through LOOCV may be less appropriate, as it leans more towards assessing imputing qualities and short range predictions than predictive performance for longer range as is usually implied by "out-of-sample" prediction. This prompts the question: What is a suitable approach moving forward?
@@ 1.3 LGOCV: Complementing LOOCV for dependent cases
Here, the group (denoted by $I_{i}$) is an index set including $i$. This configuration facilitates that the pair $(y_{i},\boldsymbol{y}_{-I_{i}})$ mimics a specified prediction task, with $\boldsymbol{y}_{-I_{i}}$ being the data subset excluding the data indexed by $I_{i}$. In a multilevel model, as depicted in Figure 1, predicting a student’s grade from an unseen class necessitates that $I_{i}$ includes $i$ and all observations from student $i$’s class. However, more complex models, such as models containing both time series and hierarchical elements, pose challenges when defining a natural prediction task. Therefore, even in complex cases, LOOCV is often applied for its simplicity—even if it leans more towards imputation or short-range prediction.
Developing a framework that evaluates a model’s longer-range prediction like the proposed LGOCV, necessitates the construction of the leave-out group $I_{i}$ for each datapoint $y_{i}$. Our approach constructs a model-based group, $I_{i}$, for each $i$ by using the prior or posterior correlation among the set of linear predictors. Though we will delve into the construction of $I_{i}$ in Section 3, an initial understanding is that $I_{i}$ comprises the data points that correspond to the linear predictors that are most informative for predicting the testing linear predictor, and thus the testing point, $y_{i}$. This set ensures that our LGOCV focuses less on short-range prediction (interpolation) and more on longer-range prediction than LOOCV. In other words, LGOCV tests the model on more difficult prediction tasks since the most influential points are removed together with the testing point, instead of some arbitrary (possibly uncorrelated) point(s). The user needs to only provide a number that indicates the "degree of the independence" between the prediction point and the rest of the data", and we compute these groups for each datapoint in an automated way. In various practical examples, we will show how this model-based procedure produces reasonable groups. Advanced spatial examples applying the proposed method are presented by [19]. For a simple time-series example, our new approach will correspond to evaluating $\pi(y_{t}|y_{1},\ldots,y_{t-k},y_{t+k},\ldots,y_{T}),$ for fixed $k>1$. This corresponds to removing a sequence of data with length $12k-$ 1, to predict the central one. As we see, this task mimics a longer-range prediction task. Our interpretation is that LGOCV quantifies the model’s ability to predict longer-range more appropriately than LOOCV, when $k>1$, and is similar to the cross-validation procedure proposed by [5] for stationary processes.
@@
There are two key challenges to address to make our proposal practical. Firstly, we must quantify the information contributed by one data point in predicting another; this is crucial for the group construction. Secondly, we face the computational task of evaluating $u_{\text{LGOCV}}$ given a set of groups. The naive computation of LGOCV by fitting models across all potential training sets and evaluating their utility against corresponding testing points is computationally infeasible, especially given the resource-demanding nature of modern statistical models. However, these challenges can be handled elegantly within the framework of latent Gaussian models (LGMs) combined with the integrated nested Laplace approximation (INLA) inference, as detailed in [20, 21, 22, 23]. Throughout this paper, we will assume that our model is an LGM. We will discuss how to integrate the automatic group construction and the fast computation of $u_{\text{LGOCV}}$ using the INLA framework. Notably, our proposed methodology has been incorporated into the R-INLA package (www.rinla.org), extending its applicability across all LGMs supported by R-INLA.
@@ 1.4 Theoretical aspects
Cross-validation (CV), particularly LOOCV, is frequently considered as an estimator of $E_{\tilde{y}}[\log\pi(\tilde{y}|\boldsymbol{y})]$ or $E_{\tilde{y},\boldsymbol{y}}[\log\pi(\tilde{y}|\boldsymbol{y})]$. The first expectation describes the generalized predictive performance given a specific training set, while the second expectation describes the generalized predictive performance averaged over different identically distributed training sets. These expectations can be evaluated when assuming the existence of the joint density $\pi_{T}(\tilde{y},\boldsymbol{y})$, representing the true data generation process. Under the assumption of exchangeability and some regularity conditions on the model, the Bernstein-Von-Mises theorem states that $\log\pi(\tilde{y}|\boldsymbol{y})$ converges to a random variable irrelevant to $\boldsymbol{y}$. Consequently, $E_{\tilde{y}}[\log\pi(\tilde{y}|\boldsymbol{y})]$ and $E_{\tilde{y},\boldsymbol{y}}[\log\pi(\tilde{y}|\boldsymbol{y})]$ become equivalent in the limit. If we further assume that $\tilde{y}$ is sampled from the same distribution as all the training data, LOOCV is an asymptotically unbiased estimator of the expectations. Commonly used information criteria, such as AIC [24], WAIC [25], are asymptotically equivalent to LOOCV in fully exchangeable cases. This type of analysis is prevalent in the literature with various settings [1, 26, 27, 28].
However, a similar analysis does not hold for dependent cases in general. Firstly, the existence of different prediction tasks means that both the model prediction, $\pi(\tilde{y}|\boldsymbol{y})$, and the true data generation process, $\pi_{T}(\tilde{y}|\boldsymbol{y})$, are not uniquely defined as discussed in Section 1.2. Secondly, the asymptotic scheme is not uniquely defined, even with a specific prediction task. For example, in a temporal model where data $\boldsymbol{y}=\{y_{1},y_{2},\dots,y_{n}\}$ is a time-series, observed at time $t_{1}<t_{2}<\dots<t_{n}$ and we denote the last time step as $T$. Several meanings of of $n\to\infty$ can be considered:
@@
These scenarios correspond to observing more future data and having higher sample rates within a time frame. As mentioned in Section 1.2, multilevel data can also have various asymptotic regimes. Thirdly, if the data generation process is not stationary, the model will not converge under certain asymptotic regimes, which differentiate $E_{\tilde{y}}[\log\pi(\tilde{y}|\boldsymbol{y})]$ from $E_{\tilde{y},\boldsymbol{y}}[\log\pi(\tilde{y}|\boldsymbol{y})]$ even in asymptotic scenarios. These points highlight that the estimand of CV is not uniquely defined in dependent cases, preventing the establishment of an asymptotic analysis framework.
From the perspective of CV, it is also inappropriate to consider it an estimator since each summand in CV should be viewed as a sample from different distributions due to the relevance of data indexes in dependent cases. For example, if we compute LOOCV in a time series. Each $y_{t}$ is sampled from a different conditional distribution $\pi_{T}(y_{t}|\boldsymbol{y}_{-t})$ and thus the average $\frac{1}{n}\sum_{t=1}^{T}\log\pi(y_{t}|\boldsymbol{y}_{-t})$ cannot be considered as an estimator in general. Therefore, it is more reasonable to view CV as a predictive measurement rather than an estimator of an expectation. This perspective allows us to interpret the proposed LGOCV as the averaged predictive performance for similar prediction tasks, created systematically by the model.
While the proposal of [16] for multilevel model demonstrates that marginal WAIC is akin to LOOCV, we note that conditional WAIC aligns with LGOCV, where a hierarchical level, such as a school, defines the groups. The h-block CV of [5] is a special case of LGOCV for a stationery model. LFOCV proposed by [9] is similar to LGOCV as shown in Section 5. The spatial buffering proposed by [11] ensures that no test data is spatially next to any training data, and is a special case of LGOCV for model with only spatial effects. LGOCV this provides a framework where no training data is placed next to the test data in terms of the entire model, and not just specific components thereof.